# Eigenvalues of Heckman-Polychronakos operators

### v1

@@ Abstract.
Heckman-Polychronakos operators form a prominent family of commuting differential–difference operators defined in terms of the Dunkl operators $\mathcal{D}_{i}$ as $\mathcal{P}_{m}=\sum_{i=1}^{N}(x_{i}\mathcal{D}_{i})^{m}$. They have been known since 1990s in connection with trigonometric Calogero-Moser-Sutherland Hamiltonian and Jack symmetric polynomials. We explicitly compute the eigenvalues of these operators for symmetric and skew-symmetric eigenfunctions, as well as partial sums of eigenvalues for general polynomial eigenfunctions.
@@ 1. Introduction
Our paper is motivated from two directions. From one side, the theory of quantum integrable systems or, more generally, spectral theory, seeks to construct families of commuting operators and understand the structure of their eigenvalues and eigenfunctions. From the other side, integrable probability often exploits operators to extract information about stochastic systems and their asymptotic behavior.
@@
We deal with a family of differential–difference operators acting on polynomials in $N$ variables $x_{1},\dots,x_{N}$ and called the Heckman-Polychronakos operators, which were first introduced in [Hec91], [Pol92] in connection to the study of (trigonometric) Calogero-Moser-Sutherland Hamiltonian. The key properties of interest for these operators were that they commute with each other and with the CMS Hamiltonian, and that Jack symmetric polynomials (whose extensive theory was developed just a couple of years before that, see [Sta89], [Mac95, Chapter VI]) are their eigenfunctions.
@@
Soon after these operators were introduced, two things happened. First, [Che91, Che94] introduced an alternative family of the Cherednik operators satisfying the same three properties, and, second, the theory was lifted up to the level of Macdonald polynomials (for whom Jack polynomials are a degeneration, cf. [Mac95, Chapter 6]). In subsequent years, Cherednik operators and their Macdonald versions were playing their roles in many developments, while Heckman-Polychronakos operators received much less attention, cf. [Cha24] for a recent historical overview of the theory. As a result, even the most basic properties of the latter operators, such as the formulas for eigenvalues, are still not well-understood by now. The goal of this paper is to fill in this gap in the literature and re-initiate the study of the Heckman-Polychronakos operators.
@@
Some relationships between Cherednik and Heckman–Polychronakos operators are discussed in [SV15, Section 5]. Both families of operators can be constructed through Newton power sums of certain “basic” operators. However, these basic operators have different properties in two theories: in Cherednik theory the basic operators commute with each other, but their definition does not agree with the natural action of $S_{N}$ on $(x_{1},\dots,x_{N})$; in Heckman–Polychronakos theory the basic operators are defined in a $S_{N}$–equivariant way, but at the cost of not commuting with each other. [^1] We emphasize that although the basic operators do not commute, but the Heckman–Polychronakos operators do commute. The two parallel possible choices of basic operators in the Macdonald $(q,t)$–theory are further discussed in [NS19], see especially Section 2.4 there.
@@
In the semiclassical limit [^2] This limit is related to degeneration of a Lie group into a Lie algebra and to an elementary transition $\lim\limits_{\varepsilon\to 0}(1+\varepsilon x)^{\varepsilon^{-1}y}=\exp(xy)$., Jack polynomials degenerate into multivariate Bessel functions (see e.g. [OO97, Section 4]) and the operators we study have a parallel degeneration to objects related to the rational (rather than trigonometric) Calogero-Moser-Sutherland Hamiltonian. Interestingly, Cherednik and Heckman–Polychronakos operators unite into a single family in this limit transition: the basic operators turn into the Dunkl operators of [Dun89] which both commute and are equivariant.
Switching to the integrable probability side, recently in [BGCG22, Xu23, GXZ24, KX24] the power sums of Dunkl operators were intensively used in the studies of $\beta$–ensembles of random matrix theory in problems related to additions of matrices and various limits of eigenvalue distributions. These random matrix distributions have natural discrete analogues [Ker00, BO05, GS15, DF16, BGG17, Mol23, DS19, GH19, Hua21, DK22, CDM23, DK24, GH24] which are intimately related to Jack polynomials. Many of the questions about matrix $\beta$–ensembles have versions for the discrete ensembles as well, and if one wants to attack them by the same techniques, then one should rely on a lifting of the power sums of Dunkl operators to the discrete setting, which would be either Cherednik or Heckman–Polychronakos operators. It is not yet clear, which of the two classes of operators would work better for the asymptotic problems, and one would like to be able to try both.
The desire to investigate the asymptotic questions about the stochastic systems from the previous paragraph was an initial motivation of our investigations of Heckman–Polychronakos operators in this paper, yet their theory turned out to be much more interesting than what we expected and, moreover, many open questions remain.
@@
Acknowledgements. We would like to thank Evgeni Dimitrov and Grigori Olshanski for very helpful discussions. V.G. was partially supported by NSF grant DMS - 2246449.
2. Heckman–Polychronakos operators
@@
Throughout the paper we fix $N=1,2,\dots$ and deal with the ring $\mathbb{C}_{N}=\mathbb{C}[x_{1},x_{2}\dots,x_{N}]$ of polynomials in $N$ variables. $\theta$ is an auxiliary parameter which enters into all our definitions: one can take it as a formal variable, or as a positive real number.
@@
For $m\geq 1$ the Heckman-Polychronakos operator $\mathcal{P}_{m}$ is a linear operator in $\mathbb{C}_{N}$:
@@
where $x_{i}$ stays for the operator of multiplication by $x_{i}$.
The operators $\mathcal{P}_{m}$ were first introduced and discussed in [Hec91], [Pol92]. The interest in them is based on three properties, whose proofs we recall in the Appendix:
@@
In contrast to another famous family of operators satisfying these three properties — power sums of the (Jack versions of) Cherednik operators [Che91, Che94] – the individual terms $(x_{i}\mathcal{D}_{i})^{m}$ do not commute. [^3] For a fruitful study of commutators of $\mathcal{D}_{i}$ and $x_{j}$ operators see [FH15]. On the other hand, the advantage of Heckman-Polychronakos operators is that their definition is simpler than the Cherednik ones and they are equivariant with respect to the symmetric group action.
@@
The appearance of $\sum_{i<j}(i,j)$ operator hints on the relevance of the representation theory of $S_{N}$ for the computations of eigenvalues and we will see that this is indeed the case.
@@ 3. Symmetric eigenfunctions
A partition of $n$ is a sequence of non-negative integers $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$, such that $\sum_{i}\lambda_{i}=n$. One can write $\lambda\vdash n$ or $|\lambda|=n$ in this situation. The length of $\lambda$, denoted $\ell(\lambda)$ is the number of non-zero parts in $\lambda$. In particular, if $\ell(\lambda)\leq N$, then $\lambda_{N+1}=0$ and we can represent $\lambda$ as $N$–tuple $(\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}\geq 0)$.
We also recall the notation $h_{m}(x_{1},\dots,x_{N})$ for complete homogeneous symmetric functions and introduce their relatives, see [Mac95, Example 19, Section I.2] for similar (but slightly different) functions:
@@
For each $1\leq r\leq N$, using the generating function for the complete homogeneous functions $h_{m}$ (see [Mac95, Chapter 1, Section 2]) we have
@@
The fact that symmetric Jack polynomials are eigenfunctions of $\mathcal{P}_{m}$ is not new: there is no doubt that the authors of [Hec91], [Pol92], or [LV96] were aware of it. However, the formulas for the eigenvalues are more elusive and we could not locate (5), (6), or (7) in the literature.
G. Olshanski pointed to us that in the special case [^4] At $\theta=1$ Jack polynomials turn into Schur functions, which are characters of the irreducible representations of the unitary group $U(N)$. $\theta=1$, (6) matches the formula for the eigenvalue of the classical Casimir operator $C_{m}$, as given in [Zhe73, (*) and Theorem 2 in Section 60].
Formula (7) is similar to the expression for eigenvalues of Nazarov-Sklyanin operators in the space of polynomials in infinitely many variables, see [NS13, Section 6].
@@ 4. Classification of eigenfunctions
Outside Jack polynomials, other eigenfunctions of $\mathcal{P}_{m}$ are much less understood, and in this section we propose a classification for them.
@@
We recall that irreducible representations of symmetric group $S_{N}$ are parameterized by partitions $\tau\vdash N$, see [Mac95, Section I.7], [OV96], [Jam06] for the reviews of the representation theory of $S_{N}$. If we denote $\mathrm{Mult}(\lambda)$ the partition of $N$ representing multiplicities of coordinates in $\lambda$ (for each $a\geq 0$ we count how many times $a$ appears in $\lambda$, and then reorder decreasingly the resulting numbers), then the summation in (8) goes over $\tau$ such that $\mathrm{Mult}(\lambda)\preceq\tau$, see, e.g. [Mac95, Remark on page 115].
Outside several special cases, for generic $\lambda$ and $\tau$ the numbers $\dim V_{\lambda;\tau}$ do not admit a closed formula, although they are of interest in algebraic combinatorics (being equivalent to Kostka numbers). It is straightforward to show that the dependence of $\dim V_{\lambda;\tau}$ on $\lambda$ is only through $\mathrm{Mult}(\lambda)$.
@@
The functions $F_{\lambda,\tau,i}$ span $\mathbb{C}_{N}$ and there are no other polynomial eigenfunctions.
@@
We do not propose any explicit form for the dependence of $F_{\lambda,\tau,i}$ on $i$. One reason is that the combination of operators $\mathcal{P}_{m}$ and the structure of $S_{N}$–representation is not sufficient to distinguish them. Indeed, for any coefficients $\alpha$ and $\beta$ and any $g\in S_{N}$, the function $\alpha\cdot F_{\lambda,\tau,i}+\beta\cdot g\circ F_{\lambda,\tau,i}$ is an eigenfunction of $\mathcal{P}_{m}$ satisfying the same three properties and with the same eigenvalue.
@@
For each $\lambda$, one can choose $\tau$ to be a one-part partition, $\tau=(N)$, corresponding to trivial representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the sum of all monomials in $V_{\lambda}$. The corresponding eigenfunction is the symmetric Jack polynomial.
For each $\lambda$ with distinct parts, one can choose $\tau$ to be a unique $N$–parts partition $\tau=1^{N}$, corresponding to the sign representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the signed sum of all monomials in $V_{\lambda}$, i.e. $\sum_{\sigma\in S_{N}}(-1)^{\sigma}\circ x_{\sigma(1)}^{\lambda_{1}}x_{\sigma(2)}^{\lambda_{2}}\cdots x_{\sigma(N)}^{\lambda_{N}}$. The corresponding eigenfunction is a skew-symmetric polynomial, which can be expressed as alternating sum of all non-symmetric Jack polynomials labeled by $N!$ compositions obtained by permuting the coordinates of $\lambda$ or, alternatively, as the symmetric Jack polynomial with shifted $\theta$ and multiplied by the $\prod_{i<j}(x_{i}-x_{j})$, see [BF97, (2.40)].
@@
For $\lambda$ with distinct parts, the space $V_{\lambda}$ is isomorphic to regular representation of $S_{N}$ of dimension $N!$. For each partition $\tau\vdash N$, $\dim V_{\lambda;\tau}$ is the square of the dimension of irreducible representation $\tau$, and this is the number of eigenfunctions $F_{\lambda,\tau,i}$.
@@
Consider a scalar product on $\mathbb{C}_{N}$, where $\overline{g}$ stays for complex conjugation:
@@
that is tautologically true by renaming $\phi_{i}\leftrightarrow\phi_{j}$ in the integral. It remains to check that $\mathbf{i}\frac{\partial}{\partial\phi_{i}}+\theta\sum_{j\neq i}\frac{1}{1-\exp(\mathbf{i}(\phi_{j}-\phi_{i}))}$ is self-adjoint. For that we record the fact that $\frac{\partial}{\partial\phi_{i}}$ derivative of the integrand in the right-hand side of (12) integrates to zero:
@@
Since operators $\mathcal{P}_{m}$ are self-adjoint by Lemma 4.7, commutative by Lemma 7.1 in the appendix, and essentially finite-dimensional (because $\mathcal{P}_{m}$ preserves the space of degree $k$ polynomials for each $k$), they have a joint eigenbasis and we only need to show that the eigenfunctions satisfy the claimed properties. Consider the space $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. By Lemma 4.6, this is an invariant space for $\mathcal{P}_{m}$. Because it is finite-dimensional and $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, this space has a complete system of eigenfunctions of $\mathcal{P}_{m}$. Inductively counting eigenfunctions, which belong to such spaces with smaller $\lambda$, we conclude that in $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ there are precisely $\dim V_{\lambda}$ eigenfunctions, which have at least one non-zero coefficient among monomials from $V_{\lambda}$: these are various $F_{\lambda,\tau,i}$ with fixed $\lambda$ and varying $\tau$ and $i$. By triangularity of Lemma 4.6, for each such function, its part spanned by the monomials from $V_{\lambda}$ is necessary an eigenfunction of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. The latter operator has a symmetric expression in terms of $x_{1},\dots,x_{N}$, and therefore it commutes with the action of $S_{N}$. Hence, by the Schur’s lemma, each $V_{\lambda;\tau}$ is an invariant subspace for $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. Therefore, there are precisely $\dim V_{\lambda;\tau}$ eigenfunctions of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$ inside $V_{\lambda;\tau}$. By triangularity, each of them has a unique extension to an eigenfunction of $\mathcal{P}_{m}$ and these are $F_{\lambda,\tau,i}$, $1\leq i\leq\dim V_{\lambda;\tau}$. ∎

### v2

@@ Abstract.
Heckman-Polychronakos operators form a prominent family of commuting differential-difference operators defined in terms of the Dunkl operators $\mathcal{D}_{i}$ as $\mathcal{P}_{m}=\sum_{i=1}^{N}(x_{i}\mathcal{D}_{i})^{m}$. They have been known since 1990s in connection with trigonometric Calogero-Moser-Sutherland Hamiltonian and Jack symmetric polynomials. We explicitly compute the eigenvalues of these operators for symmetric and skew-symmetric eigenfunctions, as well as partial sums of eigenvalues for general polynomial eigenfunctions.
@@ 1. Introduction
Our paper is motivated by two directions. On one hand, the theory of quantum integrable systems, or more generally spectral theory, seeks to construct families of commuting operators and to understand the structure of their eigenvalues and eigenfunctions. On the other hand, integrable probability often exploits such operators to extract information about stochastic systems and their asymptotic behavior.
@@
We investigate a family of differential-difference operators acting on polynomials in $N$ variables $x_{1},\dots,x_{N}$, known as the Heckman–Polychronakos operators, which were first introduced in [Hec91, Pol92] in connection with the study of the (trigonometric) Calogero–Moser–Sutherland Hamiltonian. The key properties of interest for these operators were that they commute with each other and with the Calogero-Moser-Sutherland Hamiltonian, and that Jack symmetric polynomials (whose extensive theory was developed just a couple of years before that, see [Sta89], [Mac95, Chapter VI]) are their eigenfunctions.
@@
Shortly after their introduction, two important developments occurred. First, [Che91, Che94] introduced an alternative family of Cherednik operators satisfying the same three properties. Second, the theory was lifted to the level of Macdonald polynomials, of which Jack polynomials are a degeneration (cf. [Mac95, Chapter 6]). In the following years Cherednik operators and their Macdonald generalizations played central roles in many developments, while the Heckman-Polychronakos operators received significantly less attention (cf. [Cha24] for a recent historical overview). As a result, even the most basic properties of the latter operators, such as explicit formulas for their eigenvalues, remain poorly understood. The goal of this paper is to help fill this gap in the literature and to reinitiate the study of the Heckman–Polychronakos operators.
@@
Some relationships between Cherednik and Heckman–Polychronakos operators are discussed in [SV15, Section 5]. Both families of operators can be constructed via Newton power sums of certain ”basic” operators. However, these basic operators exhibit different properties in the two theories: in Cherednik theory, the basic operators commute with each other, but their definition is not compatible with the natural action of the symmetric group $S_{N}$ on $(x_{1},\dots,x_{N})$; in Heckman–Polychronakos theory, the basic operators are defined in an $S_{N}$-equivariant way, but at the cost of losing commutativity. [^1] We emphasize that although the basic operators $x_{i}\mathcal{D}_{i}$ do not commute, the Heckman–Polychronakos operators themselves do commute; see Definition 2.2, Lemma 7.1, and (38). The two parallel choices of basic operators in Macdonald $(q,t)$ theory are further discussed in [NS19], especially in Section 2.4.
@@
In the semiclassical limit, [^2] This limit is related to the degeneration of a Lie group into a Lie algebra and to the elementary transition $\lim\limits_{\varepsilon\to 0}(1+\varepsilon x)^{\varepsilon^{-1}y}=\exp(xy)$. Jack polynomials degenerate into multivariate Bessel functions (see, e.g., [OO97, Section 4]), and the operators we study degenerate in parallel to objects related to the rational (rather than trigonometric) Calogero–Moser–Sutherland Hamiltonian. Interestingly, in this limit the Cherednik and Heckman–Polychronakos operators merge into a single family: their basic operators converge to the Dunkl operators of [Dun89], which both commute and are $S_{N}$-equivariant.
Switching to the integrable probability side, recent works [BGCG22, Xu23, GXZ24, KX24] have made intensive use of power sums of Dunkl operators in the study of $\beta$-ensembles in random matrix theory. These studies address problems such as the addition of matrices, corner cutting, and the analysis of classical random matrix ensembles. Typical questions in this area concern the asymptotic behavior of individual eigenvalues, as well as more intricate functions of eigenvalues that capture their joint distribution as the matrix size tends to infinity. For the operators to be useful in such asymptotic analyses, two features are particularly important. First, the operators should have reasonably simple combinatorial expressions, allowing their action on explicit functions to be tractable in asymptotic regimes. Second, the eigenvalues of the operators correspond to observables of the stochastic system of interest, and thus should also be reasonably explicit. The power sums of Dunkl operators satisfy both of these criteria.
The random matrix distributions discussed above have natural discrete analogues, extensively studied in [Ker00, BO05, GS15, DF16, BGG17, DS19, GH19, Hua21, DK22, CDM23, Mol23, DK24, GH24], and closely connected to Jack polynomials. Many questions posed for matrix $\beta$-ensembles have discrete counterparts, and similar techniques may be applied to analyze them. This, however, requires lifting the power sums of Dunkl operators to the discrete setting — leading naturally to either Cherednik or Heckman–Polychronakos operators. It is not yet clear which of the two classes is better suited for asymptotic analysis, and ideally, one would like to explore both.
The desire to study asymptotic properties of these stochastic systems was the original motivation behind our investigation of Heckman–Polychronakos operators. However, their theory turned out to be far richer than expected, revealing many additional structural features and open questions that remain to be explored.
The rest of the paper is organized as follows. In Section 2 we formally define the Heckman-Polychronakos operators. In Section 3 we present a formula for their eigenvalues on symmetric eigenfunctions. In Section 4 we propose a classification of general polynomial eigenfunctions according to the degrees in their leading monomials and the type of representation of the symmetric group $S_{N}$ to which these monomials belong. In Section 5 we relate partial sums of eigenvalues to values of the characters of the symmetric group and explain how this can be used for explicit computations of some of the eigenvalues. In Section 6 we compute a three variable example highlighting that in some cases the eigenvalues are much more complicated. Finally, Section 7 is an appendix, where we recall various useful properties of the Dunkl operators.
@@
Acknowledgements. We would like to thank Evgeni Dimitrov and Grigori Olshanski for very helpful discussions. We thank the referees for reading the manuscript and providing valuable suggestions. V.G. was partially supported by NSF grant DMS - 2246449.
2. Heckman-Polychronakos operators
@@
Throughout the paper we fix $N=1,2,\dots$ and use the ring $\mathbb{C}_{N}=\mathbb{C}[x_{1},x_{2}\dots,x_{N}]$ of polynomials in $N$ variables. We let $\theta$ be an auxiliary parameter which will either play the role of a formal variable or a positive number throughout the paper.
@@
For $m\geq 1$ the Heckman-Polychronakos operator $\mathcal{P}_{m}$ is a linear operator in $\mathbb{C}_{N}$ acting by
@@
where $x_{i}$ stands for the operator of multiplication by $x_{i}$.
The operators $\mathcal{P}_{m}$ were first introduced and discussed in [Hec91, Pol92]. The interest in them is based on three properties, whose proofs we recall in the Appendix:
@@
In contrast to another famous family of operators satisfying these three properties — power sums of the (Jack versions of) Cherednik operators [Che91, Che94] — the individual terms $(x_{i}\mathcal{D}_{i})^{m}$ do not commute. [^3] For a fruitful study of commutators of $\mathcal{D}_{i}$ and $x_{j}$ operators see [FH15]. On the other hand, the advantage of Heckman-Polychronakos operators is that their definition is simpler than the Cherednik ones and they are equivariant with respect to the symmetric group action.
@@
The appearance of $\sum_{i<j}(i,j)$ operator hints on the relevance of the representation theory of symmetric group $S_{N}$ for the computations of eigenvalues and we will see that this is indeed the case.
@@ 3. Symmetric eigenfunctions
A partition of $n$ is a sequence of non-negative integers $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$, such that $\sum_{i}\lambda_{i}=n$. One can write $\lambda\vdash n$ or $|\lambda|=n$ in this situation. The length of $\lambda$, denoted $\ell(\lambda)$ is the number of non-zero parts in $\lambda$. In particular, if $\ell(\lambda)\leq N$, then $\lambda_{N+1}=0$ and we can represent $\lambda$ as $N$-tuple $(\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}\geq 0)$. Sometimes exponents are used to indicate multiplicity, e.g. $(3,1^{2})=(3,1,1)$.
We also recall the notation $h_{m}(x_{1},\dots,x_{N})$ for complete homogeneous symmetric polynomials and introduce their relatives, see [Mac95, Example 19, Section I.2] for similar (but slightly different) polynomials:
@@
For each $1\leq r\leq N$, using the generating function for the complete homogeneous polynomials $h_{m}$ (see [Mac95, Chapter 1, Section 2]) we have
@@
The fact that symmetric Jack polynomials are eigenfunctions of $\mathcal{P}_{m}$ is not new: there is no doubt that the authors of [Hec91, Pol92, LV96] were aware of it. However, the formulas for the eigenvalues are more elusive and we could not locate (5), (6), or (7) in the literature.
G. Olshanski pointed out to us that in the special case [^4] At $\theta=1$ Jack polynomials turn into Schur polynomials, which are characters of the irreducible representations of the unitary group $U(N)$. $\theta=1$, (6) matches the formula for the eigenvalue of the classical Casimir operator $C_{m}$, as given in [Zhe73, (*) and Theorem 2 in Section 60].
Formula (7) is similar to the expression for eigenvalues of Nazarov-Sklyanin operators in the space of polynomials in infinitely many variables, see [NS13, Section 6], and also [MM23] for further references and discussions related to these operators.
@@ 4. Classification of eigenfunctions
Outside Jack polynomials, other eigenfunctions of $\mathcal{P}_{m}$ are much less understood. In this section, we propose a classification for them based on the degrees of their leading monomials and the isotype (equivalence class) of the representation of the symmetric group $S_{N}$ to which these monomials belong.
@@
We recall that irreducible representations of symmetric group $S_{N}$ are parameterized by partitions $\tau\vdash N$, see [Mac95, Section I.7], [OV96], or [Jam06] for the reviews of the representation theory of $S_{N}$. If we denote $\mathrm{Mult}(\lambda)$ the partition of $N$ representing multiplicities of coordinates in $\lambda$ (for each $a\geq 0$ we count how many times $a$ appears in $\lambda$, and then reorder decreasingly the resulting numbers), then the summation in (8) goes over $\tau$ such that $\mathrm{Mult}(\lambda)\preceq\tau$, see, e.g. [Mac95, Remark on page 115].
Outside several special cases, for generic $\lambda$ and $\tau$ the numbers $\dim V_{\lambda;\tau}$ do not admit a closed formula, although they are of interest in algebraic combinatorics (being equivalent to Kostka numbers, cf. [Mac95, Section I.6]). It is straightforward to show that the dependence of $\dim V_{\lambda;\tau}$ on $\lambda$ is only through $\mathrm{Mult}(\lambda)$.
@@
The polynomials $F_{\lambda,\tau,i}$ span $\mathbb{C}_{N}$ and there are no other polynomial eigenfunctions.
@@
We do not propose any explicit form for the dependence of $F_{\lambda,\tau,i}$ on $i$. One reason is that the combination of operators $\mathcal{P}_{m}$ and the structure of $S_{N}$-representation is not sufficient to distinguish them. Indeed, for any coefficients $\alpha$ and $\beta$ and any $g\in S_{N}$, the polynomial $\alpha\cdot F_{\lambda,\tau,i}+\beta\cdot g\circ F_{\lambda,\tau,i}$ is an eigenfunction of $\mathcal{P}_{m}$ satisfying the same three properties and with the same eigenvalue.
@@
For each $\lambda$, one can choose $\tau$ to be a one-part partition, $\tau=(N)$, corresponding to the trivial representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the sum of all monomials in $V_{\lambda}$. The corresponding eigenfunction is the symmetric Jack polynomial.
For each $\lambda$ with distinct parts, one can choose $\tau$ to be a unique $N$-parts partition $\tau=1^{N}$, corresponding to the sign representation of $S_{N}$. Then $V_{\lambda;\tau}$ is one dimensional space spanned by the signed sum of all monomials in $V_{\lambda}$, i.e. $\sum_{\sigma\in S_{N}}(-1)^{\sigma}\circ x_{\sigma(1)}^{\lambda_{1}}x_{\sigma(2)}^{\lambda_{2}}\cdots x_{\sigma(N)}^{\lambda_{N}}$. The corresponding eigenfunction is a skew-symmetric polynomial, which can be expressed as alternating sum of all non-symmetric Jack polynomials labeled by $N!$ compositions obtained by permuting the coordinates of $\lambda$ or, alternatively, as the symmetric Jack polynomial with shifted $\theta$ and multiplied by the $\prod_{i<j}(x_{i}-x_{j})$, see [BF97, (2.40)].
@@
For $\lambda$ with distinct parts, the space $V_{\lambda}$ is isomorphic to the regular representation of $S_{N}$ of dimension $N!$, because there is a bijection between elements of $S_{N}$ and monomials $x^{\gamma}$ with $\gamma^{+}=\lambda$. For each partition $\tau\vdash N$, $\dim V_{\lambda;\tau}$ is the square of the dimension [^5] The dimension can be computed, e.g., by the hook length formula of [FRT54]. of the irreducible representation $\tau$, and this is the number of eigenfunctions $F_{\lambda,\tau,i}$.
@@
Consider a scalar product on $\mathbb{C}_{N}$, where $\overline{g}$ stands for complex conjugation and $\theta\geq 0$:
@@
that is tautologically true by renaming $\phi_{i}\leftrightarrow\phi_{j}$ in the integral. It remains to check that $\mathbf{i}\frac{\partial}{\partial\phi_{i}}+\theta\sum_{j\neq i}\frac{1}{1-\exp(\mathbf{i}(\phi_{j}-\phi_{i}))}$ is self-adjoint. For that we note that if $F(\phi)$ is periodic and differentiable on the torus $0\leq\phi<2\pi$, then $\int_{0}^{2\pi}\frac{\partial}{\partial\phi}F\left(\phi\right)\mathrm{d}\phi=F\left(2\pi\right)-F\left(0\right)=0$. Hence, recording the fact that the $\frac{\partial}{\partial\phi_{i}}$ derivative of the integrand in the right-hand side of (12) integrates to zero we obtain:
@@
The operators $\mathcal{P}_{m}$ are self-adjoint by Lemma 4.7, commutative by Lemma 7.1 in the appendix, and essentially finite-dimensional because $\mathcal{P}_{m}$ preserves the space of degree $k$ polynomials for each $k$. Hence, they have a joint eigenbasis and we only need to show that the eigenfunctions satisfy the claimed properties. Consider the space $\bigoplus\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. By Lemma 4.6, this is an invariant space for $\mathcal{P}_{m}$. Because it is finite-dimensional and $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, this space has a complete system of eigenfunctions of $\mathcal{P}_{m}$. Inductively counting eigenfunctions, which belong to such spaces with smaller $\lambda$, we conclude that in $\bigoplus\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ there are precisely $\dim V_{\lambda}$ eigenfunctions, which have at least one non-zero coefficient among monomials from $V_{\lambda}$: these are various $F_{\lambda,\tau,i}$ with fixed $\lambda$ and varying $\tau$ and $i$. By triangularity of Lemma 4.6, for each such polynomial, its part spanned by the monomials from $V_{\lambda}$ is necessarily an eigenfunction of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. The latter operator has a symmetric expression in terms of $x_{1},\dots,x_{N}$, and therefore it commutes with the action of $S_{N}$. Hence, by the Schur’s lemma, each $V_{\lambda;\tau}$ is an invariant subspace for $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$. Therefore, there are precisely $\dim V_{\lambda;\tau}$ eigenfunctions of $(T_{1}^{m}+T_{2}^{m}+\dots+T_{N}^{m})$ inside $V_{\lambda;\tau}$. By triangularity, each of them has a unique extension to an eigenfunction of $\mathcal{P}_{m}$ and these are $F_{\lambda,\tau,i}$, $1\leq i\leq\dim V_{\lambda;\tau}$. ∎